OpenStudy (anonymous):

Solve the exponential equation. Express the solution set in terms of natural logarithms. 83x = 2.2

OpenStudy (anonymous):

exponential?

OpenStudy (anonymous):

the final answer should be a natural logarithm

OpenStudy (anonymous):

um wat is the qn? is that 83*x=2.2?

OpenStudy (anonymous):

8^(3x) = 2.2

OpenStudy (anonymous):

3xln2^3=ln2.2 x=ln2.2/9ln2